The Acmeville Metropolitan Bus Service currently charges $0.77 for an all-day ticket, and has an average of 472 riders a day. The bus company is not earning a profit, but according to their contract with the city, they cannot cut the number of buses on the road. They must therefore find a way to increase revenues.

The bus company is considering increasing the ticket price to $0.88. The marketing department's studies indicate this price increase would reduce usage to 208 riders per day.

Calculate the absolute value of the price elasticity of demand for bus tickets using the simple percentage change method. Round your answer to one decimal place.

------------------

I did (208-472) / ((208+472)/2), then the same for the price, and then divided quantity/price. I was very wrong haha.
Answer #1

Using simple percentage method :Price elasticity of demand = (Percentage change in quantity demanded/ Percentage change in price)

% change in quantity demanded = [(208-472)/472]100 = (-264/472)100 = - 55.93%

% change in price = [(0.88-0.77)/0.77]100 = (0.11/0.77)100 = 14.285%

Elasticity of demand = (-55.93 / 14.285) = -3.92

SO, absolute value of price elasticity of demand = 3.92.